OpenStudy (anonymous):

-9x+7=3x+19 Show Work.

OpenStudy (anonymous):

add 9x to both sides

OpenStudy (anonymous):

now you have 7 = 12x + 19

OpenStudy (anonymous):

subtract 19 from both sides

OpenStudy (anonymous):

you get : -12 = 12x

OpenStudy (anonymous):

divide by 12 from both sides to get -1 = x

OpenStudy (anonymous):

^ yup thats about it -9x+7=3x+19 -3 -3 -12x+7=19 -7 -7 -12x=12 (divide both sides by 12) x= -1

OpenStudy (anonymous):

^.^